OpenStudy (anonymous):

Differentiate the function with respect to x y=x^g(cosx) help me with this please !!!!

OpenStudy (anonymous):

take log on both side lny=g(cos(x))x now take derivative and use chain rule 1/y dy/dx=g(cos(x))+xg'(cos(x)) (-sin(x)